# Distributed Maximum Flow in Planar Graphs

### v3

@@ 5 Dual Distance Labeling
Section organization. First, in Section 5.1 we extend the analysis of the BDD, divided into three subsections, in each we provide the proofs of some new properties of the BDD. In particular, Section 5.1.1 focuses on analyzing the way faces get split, Section 5.1.2 focuses on the dual decomposition, and Section 5.1.3 shows how to learn that decomposition distributively. After which, we would be ready to formally define the labeling scheme in Section 5.2. Finally, in Section 5.3 we show an algorithm that constructs the labels and deduces an SSSP tree.
@@ 5.1 Extending the BDD
Since our focus is on the dual graph, we need to understand how faces of $G$ are affected by the BDD. Namely, decomposing a bag $X$ into child bags $X_{1},X_{2},\ldots$ might partition some faces of $X$ into multiple parts, where each such face-part is contained in a distinct child bag of $X$. We next discuss this in detail. Since each edge belongs to two faces of $G$, it is convenient to view an edge as two copies (called darts) in opposite directions. This is fairly common in the centralized literature on planar graphs.
@@
Darts. Every edge $e\in E(X)$ is represented by two darts $d^{-}$ and $d^{+}$ (if the graph is directed, $d^{+}$ has the same direction as $e$, and $d^{-}$ is in the opposite direction). The reverse dart is defined as $\mathsf{rev}(d^{+})=d^{-}$ and $\mathsf{rev}(d^{-})=d^{+}$. We think of the two darts as embedded one on top of the other (i.e., replacing the edges with darts does not create new faces). When we mention a path or a cycle $C$ of darts, we denote by $\mathsf{rev}(C)$ the path or cycle obtained by reversing all darts of $C$. The faces of $G$ define a partition over the set of darts (i.e., each dart is in exactly one face) such that each face of $G$ is a cycle of clockwise directed darts. See Figure 9.
@@
Figure 10: Two levels of the BDD (blue) and their corresponding levels in the dual decomposition (red). For simplicity, the dual node corresponding to the face $f_{\infty}$ of $G$ is not illustrated in $G^{*}$ (its incident edges are the ones missing an endpoint).
@@ 1. Few face-parts property:
    Each bag $X$ contains at most $O(\log n)$ face-parts. Moreover, the number of face-parts resulting from partitioning faces and face-parts of $X$ across its child bags is $O(D\log^{2}n)$.
@@ 3. Dual bags’ properties:
    Let $X^{*}$ be a non-leaf bag with child bags $X^{*}_{1},X^{*}_{2},\ldots$; Then, $X^{*}$ is equivalent to $\cup_{i}X^{*}_{i}\cup X^{*}[S_{\!{}_{X}}]$ after connecting all node-parts $f_{1},f_{2},\ldots$ in $X^{*}_{1},X^{*}_{2},\ldots$ (resp.) corresponding to the same $f\in F_{\!{}_{X}}$ with a clique and contracting it.  [^5] For an intuition, see Figure 10: merging $f_{1}^{1}$ and $f_{1}^{2}$ to one node would result with $G^{*}$ after adding the separator edges.
@@ 7. Distributed knowledge:
    Each node $f$ of $X^{*}$ has an $\tilde{O}(1)$ ID, s.t. the vertices $v$ of $X$ lying on $f$ know its ID. Moreover, $v$ knows whether $f$ corresponds to a face, a face-part or the critical face of $X$. The IDs are such that the ID of a face (face-part) $f$ of $X$ is contained in the ID of face-parts $f^{\prime}$ resulting from partitioning $f$ in child bags of $X$.
@@
    For all bags $X$, its vertices know the dual edges in $X^{*}$ (if exist) corresponding to each of their incident edges.
@@ 5.1.1 The Few Face-parts Property
We start by proving Property 9.
@@
Since the height of $\mathcal{T}$ is $O(\log n)$, it is sufficient to prove that in each bag $X$ of $\mathcal{T}$ there is at most one face of $G$ that is entirely contained in $X$ and is partitioned into face-parts in $X$’s child bags. [^6] In [26], a single hole of $X$ (the union of some $S_{\!{}_{H}}$ edges of ancestors $H$ of $X$, referred to in [26] as the boundary of $X$) might be partitioned between $X$’s child bags. In our case, since we do not care about holes, we do not count this as a face (or a face-part) that gets partitioned. This would imply that each face-part of $X$ can be associated with the unique leafmost ancestor of $X$ in which this face was whole (and thus there are only $O(\log n)$ face-parts in $X$). Note, in bags $X$ of deeper recursion levels, a face-part might consist of multiple disconnected paths, we still count it as one face-part of $X$ since all of those paths belong to the same face of $G$.
To show that indeed there is only one such face, we need to understand the cases in which faces are partitioned in the BDD of [26]. The set $S_{\!{}_{X}}$ in a bag $X\in\mathcal{T}$ is a simple (unoriented) cycle consisting of two paths in a spanning tree $T$ of $X$ and an additional edge $e_{\!{}_{X}}$. The edge $e_{\!{}_{X}}$ might not be an edge of $E(G)$, in which case, we refer to it as the virtual edge of $X$. We think of $e_{\!{}_{X}}$ as if it is temporarily added to $X$, in order to define its child bags, and then removed. The cycle $S_{\!{}_{X}}$ is a balanced cycle separator, i.e. a cycle that contains a constant factor of $V(G)$ in each of its interior and exterior.
@@
In the root of $\mathcal{T}$, we have the graph $G$ and the claim trivially holds. Assume the claim holds for level $\ell-1$ of the BDD, we prove for level $\ell$. Recall that for a bag $X$, every edge $e$ of $X$ either belongs to one child bag of $X$ or to two child bags of $X$ (if $e\in S_{\!{}_{X}}$). If $e$ is not an edge of $S_{\!{}_{X}}$, then $e$ belongs to one child bag $X_{i}$ and its (one or two) darts in $X$ belong to $X_{i}$ (as both of $e$’s endpoints are in $X_{i}$). If $e$ is an edge of $S_{\!{}_{X}}$ that is contained only in one child bag, then by the BDD algorithm $e$ must be an edge of $S_{\!{}_{H}}$ for some ancestor $H\neq X$ of $X$. Thus, $e$’s only dart is also contained in the same child bag that contains $e$. Finally, if $e$ is an edge of $S_{\!{}_{X}}$ that is contained in two child bags of $S_{\!{}_{X}}$, then $e$ is in $X$ and does not participate in $S_{\!{}_{H}}$ of any ancestor $H$, then, $e$ has both of its darts in $X$. Thus, $e$ belongs to two child bags of $X$, in particular, one child bag is the internal (to $S_{\!{}_{X}}$) and the other is external, each would contain one dart of $e$ as each dart of $e$ is in a face (not a hole), and in this case, $S_{\!{}_{X}}$ separates those two faces (face-parts), each to a child bag. That way, all darts that correspond to the same face go to the same child bag of $X$. ∎
@@ 5.1.2 Dual Bags and Their Properties
We are finally prepared to prove the additional properties 1011, and 12 of the BDD, starting with a proof for Property 10.
@@
(if) Note, the dual edges of $S_{\!{}_{X}}$ contained in $X^{*}$ are not contained in any child bag of it as they lie on holes in them (Lemma 5.5), thus, their endpoints are in $F_{\!{}_{X}}$. In addition, the nodes of $X^{*}$ that get divided correspond to the critical face and a subset of face-parts in $X$ (Corollary 5.6), thus, they have at least two distinct edges, each in a distinct child bag of $X^{*}$, hence, are in $F_{\!{}_{X}}$.
@@
Let $K_{f}$ be the clique on the set of node-parts across child bags of $X^{*}$ corresponding to $f\in F_{\!{}_{X}}$. Denote $M:=\left(\cup_{i}X^{*}_{i}\right)\bigcup\left(X^{*}[S_{\!{}_{X}}]\right)\bigcup\left(\cup_{f\in F_{\!{}_{X}}}K_{f}\right)$. We start by showing that $M$ is very close to $X^{*}$, then show that the adjustment (contraction) suggested in the property on $M$ indeed yields $X^{*}$.
First, note that all edges of $X^{*}$ are contained in $M$, as any non-$S_{\!{}_{X}}$ edge is contained in a child bag (Lemma 5.5) and $S_{\!{}_{X}}$ edges that are contained in $X^{*}$ are added to $M$ by definition. In addition, each node of $X^{*}$, except for some nodes of $F_{\!{}_{X}}$, is present in $M$, as each node that is in some $X^{*}_{i}$ is contained in $M$ by definition, and any node of $X^{*}$ that is not in any $X^{*}_{i}$ must be an $F_{\!{}_{X}}$ node (by Property 11).
Now, we show that the contraction produces $F_{\!{}_{X}}$ nodes, connects them correctly and removes only non-$X^{*}$ edges from $M$ (the lemma follows). Consider the graph $M^{\prime}$, given by $M$ after contracting $K_{f}$ for all $f\in F_{\!{}_{X}}$.
@@
    Consider a node $f$ that was partitioned into node-parts $f_{1},f_{2},\ldots$, such that each $f_{i}$ is in a distinct child bag of $X^{*}$. We want to show that contracting all $f_{i}$ into one node, results with one node that is incident to all $f$’s $X^{*}$ neighbors in $M^{\prime}$, this node is identified with $f$. That follows from the following claims (1) Each edge of $f$ in $X$ goes to one face-part of $f$ in the child bag that contains it (by definition of face-parts). (2) The only edges without a dual in $X^{*}$ are edges of $S_{\!{}_{H}}$ for some ancestor $H$ of $X$ (Lemma 5.5). Thus, if a primal edge does not have a dual in any of $X^{*}$’s child bags, it is either an edge of $S_{\!{}_{H}}\neq S_{\!{}_{X}}$, in which case it is not in $X^{*}$ either (the claim does not fail), or it is in $S_{\!{}_{X}}$, thus, it is added to $M$ by definition. Thus, $f$’s incident edges in $X^{*}$ are partitioned among $f_{i}$’s in $M$. Moreover, by the previous item, the contraction affects none of these edges and leaves only them incident to the resulting node, which we identify with $f$. ∎
@@ 5.1.3 Distributed Knowledge Properties
To implement our algorithm, each vertex should know the IDs of all faces (face-parts) that contain it in every bag $X\in\mathcal{T}$ (i.e., the list of nodes $f\in X^{*}$ it participates in). Here we show that the distributed knowledge properties (13 and 14) of the BDD hold, which concludes the proof of Theorem 5.2. As before, it will be convenient to think of an edge as having two copies (one for each face containing it). To simplify, we refer to these as edge copies (rather than darts) since there is no need to maintain the dart orientation. Every vertex incident to an edge $e=(u,v)$ considers itself incident to two copies $e_{1},e_{2}$ of $e$, such that, $e_{1}$ ($e_{2}$) is the first (second) copy of $e$ relative to $v$ according to a clockwise order (for the other endpoint $u$, $e_{2}$ is its first copy).
We aim to assign each face (face-part) in any bag $X$ a unique $\tilde{O}(1)$ identifier and learn for each vertex $v\in X$, the set of faces (face-parts) containing it and its adjacent edge copies. This is done by keeping track of $G$’s edge copies along the decomposition. We start with (1) a local procedure that allows the endpoints of each edge copy $e$ to learn all the bags $X$ that contain $e$, then, (2) the endpoints of each edge copy $e$ learn the face of $G$ that $e$ participates in, and (3) we show how to extend this for the next level of $\mathcal{T}$. For a proof see Appendix C. This allows us to (distributively) learn the nodes of $X^{*}$, after which, we have the necessary information to learn the edges of $X^{*}$.
@@
In a single round, each vertex $v\in G$ can learn, for each incident edge copy, a list of the bags $X\in\mathcal{T}$ that contain it.
@@
There is a $\tilde{O}(D)$-round distributed algorithm that assigns unique IDs to all faces of $G$. By the end of the algorithm, each vertex $v\in G$ learns a list (of IDs) of faces that contain it. Moreover, $v$ learns for each incident copy $e^{\prime}$ of an edge $e$ in $G$ the face ID of the face of $G$ that contains $e^{\prime}$.
@@
There is a $\tilde{O}(D)$-round distributed algorithm that assigns unique $\tilde{O}(1)$-bit IDs to all faces and face-parts in all bags $X\in\mathcal{T}$. By the end of the algorithm, each vertex $v\in G$ learns a list (of IDs) of the faces and face-parts it lies on for each bag $X$ that contains it. Moreover, $v$ learns for each incident copy of an edge $e^{\prime}$ the (ID of) the face or face-part that $e^{\prime}$ participates in for each bag $X$ that contains $e^{\prime}$. The ID of a face (face-part) $f$ of $X$ is contained in the ID of face-parts $f^{\prime}$ resulting from partitioning $f$ in child bags of $X$.

### v4

@@ 5 Dual Distance Labeling
Section organization. First, in Section 5.1 we extend the analysis of the BDD, divided into three subsections, in each we provide the proofs of some new properties of the BDD. In particular, Section 5.1.1 focuses on analyzing the way faces get split, Section 5.1.2 focuses on the dual decomposition, and Section 5.1.3 shows how to learn that decomposition distributively. After which, we would be ready to formally define the labeling scheme in Section 5.2. In Section 5.3 we show an algorithm that constructs the labels, and finally in Section 5.4 we show how to learn an SSSP tree.
@@ 5.1 Extending the BDD
Since our focus is on the dual graph, we need to understand how faces of $G$ are affected by the BDD. Namely, decomposing a bag $X$ into child bags $X_{1},X_{2},\ldots$ might partition some faces of $X$ into multiple parts, where each such face-part is contained in a distinct child bag of $X$. We next discuss this in detail. Since each edge belongs to two faces of $G$, it is convenient to view an edge as two copies (called darts) in opposite directions. This is fairly common in the centralized literature on planar graphs [26].
@@
Darts. Each edge $e\in E(X)$ is represented by two darts $d^{-}$ and $d^{+}$ (if the graph is directed, $d^{+}$ has the same direction as $e$, and $d^{-}$ is in the opposite direction). The reverse dart is defined as $\mathsf{rev}(d^{+})=d^{-}$ and $\mathsf{rev}(d^{-})=d^{+}$. We think of the two darts as embedded one on top of the other (i.e., replacing the edges with darts does not create new faces). When we mention a path or a cycle $C$ of darts, we denote by $\mathsf{rev}(C)$ the path or cycle obtained by reversing all darts of $C$. The faces of $G$ define a partition over the set of darts (i.e., each dart is in exactly one face) such that each face of $G$ is a clockwise cycle of darts [26]. See Figure 10.
@@
Figure 11: Two levels of the BDD (blue) and their corresponding levels in the dual decomposition (red). For simplicity, the figure is undirected, in addition, the dual node corresponding to the face $f_{\infty}$ of $G$ is not illustrated in $G^{*}$ (its incident edges are the ones missing an endpoint).
@@ 1. Few face-parts property:
    Each bag $X$ contains $O(\log n)$ face-parts. Moreover, the number of face-parts resulting from partitioning faces and face-parts of $X$ across its child bags is $O(D\log^{2}n)$.
@@ 3. Dual bags’ properties:
    Let $X^{*}$ be a non-leaf bag with child bags $X^{*}_{1},X^{*}_{2},\ldots$; Then, $X^{*}$ is equivalent to $\cup_{i}X^{*}_{i}\cup X^{*}[S_{\!{}_{X}}]$  [^9] We abuse notation, even though $S_{\!{}_{X}}$ is defined as a set of vertices of $X$, we denote by $X^{*}[S_{\!{}_{X}}]$ the induced subgraph of $X^{*}$ on dual edges to (the edges) of $S_{\!{}_{X}}$. after connecting all node-parts $f_{1},f_{2},\ldots$ in $X^{*}_{1},X^{*}_{2},\ldots$ (resp.) corresponding to the same $f\in F_{\!{}_{X}}$ with a clique and contracting it.  [^10] For an intuition, see Figure 11: merging $f_{1}^{1}$ and $f_{1}^{2}$ to one node would result with $G^{*}$ after adding the separator edges.
@@ 7. Distributed knowledge:
    Each node $f$ of $X^{*}$ has an $\tilde{O}(1)$-bit ID, s.t. each vertex $v$ of $X$ lying on $f$ knows its ID. Moreover, $v$ knows whether $f$ corresponds to a face, a face-part or the critical face of $G$.
@@
    For each bag $X$, and vertices $v$ of $X$, $v$ knows the dual edges in $X^{*}$ (if exist) corresponding to each of $v$’s incident edges.
@@ 5.1.1 The Few Face-parts Property
First, we prove the following lemma. Then, we derive a corollary concluding the proof of Property 9.
@@
Since the height of $\mathcal{T}$ is $O(\log n)$, it is sufficient to prove that in each bag $X$ of $\mathcal{T}$ there is at most one face of $G$ that is entirely contained in $X$ and is partitioned into face-parts in $X$’s child bags. [^11] In [27], a single hole of $X$ (the union of some $S_{\!{}_{H}}$ edges of ancestors $H$ of $X$, referred to in [27] as the boundary of $X$) might be partitioned between $X$’s child bags. In our case, since we do not care about holes, we do not count this as a face (or a face-part) that gets partitioned. This would imply that each face-part of $X$ can be associated with the unique leafmost ancestor of $X$ in which this face was whole (and thus there are only $O(\log n)$ face-parts in $X$). Note, in bags $X$ of deeper recursion levels, a face-part might consist of multiple disconnected paths. However, we still count it as one face-part of $X$ since all of those paths belong to the same face of $G$.
To show that indeed there is only one such face, we need to understand the cases in which faces are partitioned in the BDD of [27]. The set $S_{\!{}_{X}}$ in a bag $X\in\mathcal{T}$ is a simple (unoriented) cycle consisting of two paths in a spanning tree $T$ of $X$ and an additional edge $e_{\!{}_{X}}$. The edge $e_{\!{}_{X}}$ might not be an edge of $E(G)$, in which case, we refer to it as the virtual edge of $X$. We think of $e_{\!{}_{X}}$ as if it is temporarily added to $X$, in order to define its child bags, and then removed. The cycle $S_{\!{}_{X}}$ is a balanced cycle separator, i.e. a cycle that contains a constant factor of $V(G)$ in each of its strict interior and exterior [^12] Formally, the interior (resp. exterior) of $S_{\!{}_{X}}$ is defined as the subgraph that is enclosed by the cw (resp. ccw) cycle of its darts. Those definitions are standard in the literature of planar graphs. See [26]..
@@
In the root of $\mathcal{T}$, we have the graph $G$ and the claim trivially holds. Assume the claim holds for level $\ell-1$ of the BDD, we prove for level $\ell$. Recall that for a bag $X$, every edge $e$ of $X$ either belongs to one child bag of $X$ or to two child bags of $X$ (if $e\in S_{\!{}_{X}}$). If $e$ is not an edge of $S_{\!{}_{X}}$, then $e$ belongs to one child bag $X_{i}$ and its (one or two) darts in $X$ belong to $X_{i}$ (as both of $e$’s endpoints are in $X_{i}$). If $e$ is an edge of $S_{\!{}_{X}}$ that is contained only in one child bag, then by the BDD algorithm $e$ must be an edge of $S_{\!{}_{H}}$ for some ancestor $H\neq X$ of $X$. Thus, $e$’s only dart is also contained in the same child bag that contains $e$. Finally, if $e$ is an edge of $S_{\!{}_{X}}$ that is contained in two child bags of $S_{\!{}_{X}}$, then $e$ is in $X$ and does not participate in $S_{\!{}_{H}}$ of any ancestor $H$, then, $e$ has both of its darts in $X$. Thus, $e$ belongs to two child bags of $X$, in particular, one child bag is the internal (to $S_{\!{}_{X}}$) and the other is external, each would contain one dart of $e$ as each dart of $e$ is in a face (not a hole), and in this case, $S_{\!{}_{X}}$ separates those two faces (face-parts), each to a child bag. ∎
@@ 5.1.2 Dual Bags and Their Properties
We are finally prepared to prove the additional properties 10, 11, and 12 of the BDD, starting with a proof for Property 10.
@@
(if) Note, the dual edges of $S_{\!{}_{X}}$ contained in $X^{*}$ are not contained in any child bag of it as they lie on holes in them (Lemma 5.5), thus, their endpoints are in $F_{\!{}_{X}}$. In addition, the nodes of $X^{*}$ that get split correspond to the critical face and a subset of face-parts in $X$ (Corollary 5.6), thus, they have at least two distinct edges, each in a distinct child bag of $X^{*}$, hence, are in $F_{\!{}_{X}}$.
@@
Let $K_{f}$ be the clique on the set of node-parts across child bags of $X^{*}$ corresponding to $f\in F_{\!{}_{X}}$. Denote $M:=\left(\cup_{i}X^{*}_{i}\right)\bigcup\left(X^{*}[S_{\!{}_{X}}]\right)\bigcup\left(\cup_{f\in F_{\!{}_{X}}}K_{f}\right)$. We start by showing that $M$ is very ”close” to $X^{*}$, then show that the contraction suggested in the property indeed yields $X^{*}$.
First, note that all edges of $X^{*}$ are contained in $M$, as any non-$S_{\!{}_{X}}$ edge is contained in a child bag (Lemma 5.5) and $S_{\!{}_{X}}$ edges that are contained in $X^{*}$ are added to $M$ by definition. In addition, each node of $X^{*}$ is present in $M$, except for some nodes of $F_{\!{}_{X}}$, as each node that is in some $X^{*}_{i}$ is contained in $M$ by definition, and any node of $X^{*}$ that is not in any $X^{*}_{i}$ must be an $F_{\!{}_{X}}$ node (by Property 11).
Now, we show that the contraction produces $F_{\!{}_{X}}$ nodes, connects them correctly and removes only non-$X^{*}$ edges from $M$ (the lemma follows). Consider the graph $M^{\prime}$, given by $M$ after contracting $K_{f}$ for all $f\in F_{\!{}_{X}}$. We claim that $M^{\prime}=X^{*}$.
@@
    Consider a node $f$ that was partitioned into node-parts $f_{1},f_{2},\ldots$, such that each $f_{i}$ is in a distinct child bag of $X^{*}$. We want to show that contracting all $f_{i}$ into one node, results with one node in $M^{\prime}$ that is incident to the same neighbors of $f$ in $X^{*}$, this node of $M^{\prime}$ is identified with $f$. That follows from the following claims (1) Each edge of $f$ in $X$ goes to one face-part of $f$ in the child bag that contains it (by definition of face-parts). (2) The only edges without a dual in $X^{*}$ are edges of $S_{\!{}_{H}}$ for some ancestor $H$ of $X$ (Lemma 5.5). Thus, if a primal edge does not have a dual in any of $X^{*}$’s child bags, it is either an edge of $S_{\!{}_{H}}\neq S_{\!{}_{X}}$, in which case it is not in $X^{*}$ either (the claim does not fail), or it is in $S_{\!{}_{X}}$, thus, it is added to $M$ by definition. Thus, $f$’s incident edges in $X^{*}$ are partitioned among $f_{i}$’s in $M$. Moreover, by the previous item, the contraction affects none of these edges and leaves only them incident to the resulting node, which we identify with $f$ in $X^{*}$. ∎
@@ 5.1.3 Distributed Knowledge Properties
To implement our algorithm, each vertex should know the IDs of all faces (face-parts) that contain it in every bag $X\in\mathcal{T}$ (i.e., the list of nodes $f\in X^{*}$ it participates in). Here we show that the distributed knowledge properties (13 and 14) of the BDD hold, which concludes the proof of Theorem 5.2. As before, it will be convenient to think of an edge as having two darts. That is, every vertex incident to an edge $e=(u,v)$ considers itself incident to two darts $d^{+},d^{-}$ of $e$.
Even though we think of the two darts of an edge as embedded one on top of the other, it would still be useful to maintain a local order for each vertex over its incident darts, extending its local ordering of edges: In the local ordering of a vertex $v$ of its darts, a dart $d^{\prime}$ of an edge $e^{\prime}$ precedes a dart $d$ of an edge $e$ iff $e^{\prime}$ precedes $e$ in the local ordering of $v$ for its edges. To define well the local ordering of $v$ for its darts, it remains to define the order between each two darts that belong to the same edge. I.e., a vertex $v$ incident to an edge $e=(u,v)$ of darts $d^{+},d^{-}$ orders $d^{+},d^{-}$ in its local ordering, s.t. the dart that inters it precedes the dart leaving it. Note, the dart $d^{+}$ inters $v$ and leaves $u$. Symmetrically, the dart $d^{-}$ inters $u$ and leaves $v$. Meaning, this ordering is consistent (depends only on the direction of $e$).
We aim to assign each face (face-part) in any bag $X$ a unique $\tilde{O}(1)$ identifier and learn for each vertex $v\in X$, the set of faces (face-parts) containing it and its adjacent darts. This is done by keeping track of $G$’s darts along the decomposition. We start with (1) a local procedure that allows the endpoints of each dart $d$ to learn all the bags $X$ that contain $d$, then, (2) the endpoints of each dart $d$ learn the face of $G$ that $d$ participates in, and (3) we show how to extend this for the next level of $\mathcal{T}$. This allows us to (distributively) learn the nodes of $X^{*}$, after which, we have the necessary information to learn the edges of $X^{*}$.
@@
In a single round, each vertex $v\in G$ learns for each incident dart, a list of the bags $X\in\mathcal{T}$ that contain it.
Proof.
We wish to learn, for each dart, the ID of the single bag (by Lemma 5.5) in each level of $\mathcal{T}$ that contains it. For the root $G$ of $\mathcal{T}$, all vertices know that all their incident darts belong to $G$. For a general bag $X$, we assume each vertex $v\in X$ knows which of its incident darts belong to $X$, and we show how to learn this for every child bag of $X$ that contains $v$. If $v$ is in only one child bag $X_{i}$ of $X$ ($v$ knows it is in $X_{i}$), then $v$’s incident edges in $X_{i}$ are the same as in $X$. Otherwise, $v$ must be a vertex of $S_{\!{}_{X}}$ (and by the BDD properties, $v$ knows this). For each child bag $X_{i}$ that contains $v$, $v$ knows two incident edges $a,b$ such that: (1) $a$ precedes $b$ in $v$’s cw local ordering of incident edges, and (2) $v$’s incident edges in $X_{i}$ are exactly $a,b$ and all edges $I_{v}(X_{i})$ strictly between $a$ and $b$ in $v$’s local ordering of incident edges. Thus, each child bag that contains $v$ is identified with a pair of ordered edges incident to $v$. Observe that one of those edges might be the virtual edge $e_{\!{}_{X}}$ (if $v$ is an endpoint of $e_{\!{}_{X}}$). Though $e_{\!{}_{X}}$ is not an actual edge of $X$, it (and its embedding) is known to its endpoints by the BDD construction. To conclude, for any dart $d$ of an edge $e$ incident to $v$, $v$ can learn if that dart is in $X_{i}$ by checking if $d$ is the dart of $a$ leaving $v$, or $d$ is the dart of $b$ that inters $v$, or $e\in I_{v}(X_{i})$. ∎
We next show how, for all bags $X$, every vertex $v\in X$ learns the IDs of faces (face-parts) that contain it, and for each of its darts $d$ in $X$ the ID of the face (face-part) in $X$ containing $d$. We begin by showing this for $X=G$ and then for a general bag $X$.
@@
There is a $\tilde{O}(D)$-round distributed algorithm that assigns unique IDs to all faces of $G$. By the end of the algorithm, each vertex $v\in G$ learns a list (of IDs) of faces that contain it. Moreover, $v$ learns for each incident dart $d$ of an incident edge $e$ in $G$ the face ID of the face of $G$ that contains $d$.
Proof.
We first assign unique $O(\log n)$-bit IDs to the faces of the graph $G$ and let all vertices learn the IDs of all the faces of $G$ that contain them. We can do so in $\tilde{O}(D)$ rounds by using Property 4 of the face-disjoint graph (Section 3.1). Once a vertex $v$ knows all faces containing it, it remains only to assign those faces to the darts incident to $v$ (which is done locally).
Recall, each edge $e$ of $G$ has two darts $d^{+},d^{-}$ in the set of edges $E_{R}$ of $\hat{G}$, and each vertex $v$ of $G$ has $\deg(v)$ many copies $v_{i}$ in $\hat{G}\setminus V_{S}$. Moreover, $\hat{G}[E_{R}]$ is a collection of disjoint cycles, each corresponding to a face of $G$. Let $e=(u,v)\in E(G)$ be the $i$’th and $j$’th edge in $u$’s and $v$’s local embedding (in $G$), respectively. Then, the copies $u_{i+1},v_{j}$ are connected by an edge in $E_{R}$ identified with the dart that inters $v$ and leaves $u$ of $e$ ($d^{+}$ since $e$ is directed from $u$ to $v$), and the copies $u_{i},v_{j+1}$ are connected an edge in $E_{R}$ identified with the dart that inters $u$ and leaves $v$ of $e$ ($d^{-}$ in this case). Note, this order of darts is the same order that we keep in other algorithms (lemmas) for consistency. We assign each of $d^{+}$ and $d^{-}$ the component (face) ID that contains them in $\hat{G}[E_{R}]$. Since $u,v$ in $G$ simulate all their darts in $\hat{G}$, each of them knows for each of $d^{+},d^{-}$ the connected component ID in $\hat{G}[E_{R}]$ it participates in. This is the ID of the face that corresponds to that component. Thus, for each of $d^{+},d^{-}$, both $u$ and $v$ learn the correspondence to faces of $G$. ∎
@@
There is a $\tilde{O}(D)$-round distributed algorithm that assigns unique $\tilde{O}(1)$-bit IDs to all faces and face-parts in all bags $X\in\mathcal{T}$. By the end of the algorithm, each vertex $v\in G$ learns a list (of IDs) of the faces and face-parts it lies on for each bag $X$ that contains it. Moreover, $v$ learns for each incident dart $d$ the (ID of) the face or face-part that $d$ participates in for each bag $X$ that contains $d$. The ID of a face (face-part) $f$ of $X$ is contained in the ID of face-parts $f^{\prime}$ resulting from partitioning $f$ in child bags of $X$.